OpenStudy (anonymous):

The radius of the Earth is apprximately 3,960 miles, while the radius of the Moon is approximately 1,080 miles. How many times larger is the surface area of the Earth than the surface area of the Moon?

OpenStudy (cwrw238):

3960^2 / 1080^2

OpenStudy (anonymous):

whats the answer?

OpenStudy (cwrw238):

you can use our calculator

OpenStudy (anonymous):

how do i get that on here?

OpenStudy (cwrw238):

your computer should have a calculator

OpenStudy (anonymous):

Find the volume of a sphere with a diameter of 25 cm. Approximate pie as 3.14 and round your answer to the nearest hundredth.

OpenStudy (anonymous):

help me with this one too?

OpenStudy (cwrw238):

volume of a sphere is (4/3) * pi * r^3 plug in pi = 3.14 and r = 12.5 (r = half of the diameter= 1/2 * 25 = 12.5)

OpenStudy (anonymous):

can you give me the answer? please

OpenStudy (cwrw238):

no - its easy to plug into into the calculator

OpenStudy (anonymous):

1471.88 is that the answer?

OpenStudy (anonymous):

????????????????